Solve 2- square root of x+1= -5

1
Expert's answer
2021-05-03T05:54:25-0400

"Solution. \newline
2-\sqrt{x+1}=-5; \newline
-\sqrt{x+1}=-7; |*(-1) \newline
\sqrt{x+1}=7; \newline
x+1>0; \newline
x>-1; \newline
\sqrt{x+1}=7; |^2 \newline
x+1=49; \newline
x=48. \newline
Answer: x=48."